### v2

@@ Abstract
The set of relevant cuts in a graph is the union of all minimum weight bases of the cut space. We show that a cut is relevant if and only if it is the a minimum weight cut between two distinct vertices. Moreover, we give a characterization in terms of Picard-Queyranne DAGs that can be used to accelerate the enumeration of the relevant cuts.
@@ A Short Note on Relevant Cuts
Nico DomschkeBioinformatics Group, Department of Computer Science, Leipzig University, Härtelstraße 16–18, D-04107 Leipzig, Germanynico@bioinf.uni-leipzig.de
@@
Richard GolnikBioinformatics Group, Department of Computer Science, Leipzig University, Härtelstraße 16–18, D-04107 Leipzig, Germanythomas@bioinf.uni-leipzig.de
@@ 1 Introduction
Cuts in graphs are closely related to cycles. The cycle space and the cut space form orthogonal subspaces of the $|E|$-dimensional vector space over GF(2) spanned the edges of $G$. Nevertheless, graphs can contain edge sets called bi-cycles that are both cycles and cuts [24]. The cycle space, and in particular bases of the cycle space comprising short cycles, play an important role in cheminformatics [8, 15, 18] and structural biology [17, 16]. In particular, minimum cycle bases composed of cycles with minimal total length or edge weight, and their union, known as the set of relevant cycles [27], are of use [3].
@@
Minimum cut bases in undirected graphs are studied in detail in [5], their counterparts in directed graphs were analyzed in [13, 14], highlighting the special role of Gomory-Hu trees. These are (not necessarily spanning) trees on $V(G)$ with edge weights $\lambda_{e}$ such that for each pair $s,t\in V(G)$, the minimum weight $\lambda_{e}$ along the path between $s$ and $t$ in $T$ equals the weight of a minimum $s,t$-cut [9]. Since Gomory-Hu trees exist for all graphs, certain sets of $n-1$ distinct minimal cuts already provide a minimal cut for any pair of distinct vertices. As shown in [5, 13], the cuts defined by a Gomory-Hu tree form minimal cut (co-cycle) bases. Moreover, they are weakly fundamental, i.e., their cuts can be ordered in such a way that every cut contains an edge that is not present in any of its predecessors.
@@ 2 Preliminaries
Throughout, $G=(V,E,w)$ is an undirected connected graph with edge weights $w:E\to\mathbb{R}$. In the context of chemical chemical application, integer weight corresponding to the order of a chemical bond will be most relevant. A cut is a proper bipartition $(V^{\prime},V^{\prime\prime})$ of the vertex set $V(G)$, i.e., $V^{\prime}\neq\emptyset$, $V^{\prime\prime}\neq\emptyset$, $V^{\prime}\cap V^{\prime\prime}=\emptyset$, and $V^{\prime}\cup V^{\prime\prime}=V$. A cutset $D\coloneqq\operatorname{cut}(V^{\prime},V^{\prime\prime})=\{xy\in E|x\in V^{\prime},y\in V^{\prime\prime}\}$. The weight of a cutset is $|D|\coloneqq\sum_{xy\in D}w(xy)$. In the following we will usually write $\partial V^{\prime}=\partial V^{\prime\prime}=\operatorname{cut}(V^{\prime},V^{\prime\prime})=\operatorname{cut}(V^{\prime\prime},V^{\prime})$. In particular, the edges incident to each vertex $x\in V$ form a cutset $\partial\{x\}$. A cutset $D\in\mathfrak{D}$ that does not include another cutset $D^{\prime}$ of $G$ as a proper subset is called a bond of $G$. Clearly, each $\partial\{x\}$ is bond. Let $T$ be a spanning tree of $G$, and $e$ an edge of $T$. Removal of $e$ decomposed $T$ into exactly two connected components with vertex sets $V_{e}^{\prime}$ and $V_{e}^{\prime\prime}$. By construction the induced subgraphs $G[V_{e}^{\prime}]$ and $G[V_{e}^{\prime\prime}]$ are connected. The cut $\partial V_{e}^{\prime}$ is also a bond.
@@
For connected graphs, $\mathfrak{D}$ has dimension $n-1$, where $n=|V|$ is the number of vertices in $G$. In particular, each of the sets $\mathcal{D}_{x_{0}}\coloneqq\{\partial\{x\}|x\in V\setminus\{x_{0}\}\}$ for any $x_{0}\in V$, and each of the sets $\mathcal{D}_{T}\coloneqq\{\partial V_{e}^{\prime}|e\in E(T)|\}$ for any spanning tree $T$ of $G$ is basis of the cut space, see e.g. [4].
A cut $(V^{\prime},\bar{V}^{\prime})$ is an $s,t$-cut if $s\in V^{\prime}$ and $t\in\bar{V}^{\prime}$. Clearly, the cut space $\mathfrak{D}$ contains an $s,t$-cut for every pair $s\neq t$ of distinct vertices. Below, we will make use of the fact that this is also true for every basis of $\mathfrak{D}$, see e.g. also [5]. For completeness we include a short proof.
@@
Let $\mathcal{D}=\{\partial U_{i}|i=1,\dots,n-1\}$ be a cut-basis, and set $Z\coloneqq\bigcup_{i}U_{i}$. Then any cut that can be obtained as $\oplus$-sum of a subset of $\partial U_{i}$ does not contain any edge with both endpoints in $Z$ and thus must be of the form $(W^{\prime},\bar{W}^{\prime})$ with $Z\subseteq W^{\prime}$ or $Z\subseteq\bar{W}^{\prime}$. Assume $s,t\in Z$. Then the cut space generated by $\mathcal{D}$ does not contain an $s,t$-cut. This is impossible, however, since $\partial\{s\}$ is an $s,t$-cut is contained in the cut space, contradicting the assumption that $\mathcal{D}$ is a basis.
@@
    for every $st\coloneqq e\in F$, the bipartition of $V$ induced by removing $e$ from $F$ defines a cut of $G$, which we denote by $(V_{e},\bar{V}_{e})$;
@@
    for every $st\coloneqq e\in F$, the weight $\lambda_{e}$ of $e$ is the weight of a minimum $s,t$-cut;
@@
    for all $u,v\in V$, the weight of a minimum $u,v$-cut in $G$ equals the minimum weight $\lambda_{e}$ along the unique path $P_{uv}$ connecting $u$ and $v$ in $T^{*}$.
@@
[9] Every weighted graph $G=(V,E,w)$ has a Gomory Hu tree $T^{*}=(V,F,\lambda^{*})$.
@@
[5] The cut-sets $\partial{V_{e}}$ induced by the edges $e\in F$ of a Gomory-Hu tree $T^{*}$ for $G$ form a minimum cut basis of $G$.
@@ 3 Relevant Cuts
The notion of relevant cycles was and studied for cycles bases of undirected graphs in [27]. Interestingly, the notion was proposed already in 1971 by Morris Plotkin in the context of chemical ring perception [20]. More generally, the concept of relevant elements pertains to matroids with a weight function in general and reads as follows for the special case of cut spaces:
@@
The following simple characerization of relevant elements holds for matroids in general [27]:
@@
Since $\mathfrak{D}$ is a vector space it is also matroid. Suppose $D$ is contained in the minimum cut basis $\mathcal{D}$. If $D$ can be written as the $\oplus$-sum of shorter cuts $D_{1}$ through $D_{k}$, it could be exchanged with one of these cuts, contradicting minimality of $\mathcal{D}$. Now suppose that $D$ cannot be written as a sum shorter cuts. Then $D=D_{1}\oplus\dots\oplus D_{k}$ and there is at least one cut $D_{i}$ with $|D_{i}|\geq|D|$. Now $D_{i}=D_{1}\oplus\dots\oplus D_{i-1}\oplus D\oplus D_{i+1}\oplus\dots\oplus D_{k}$ and thus $\mathcal{D}^{\prime}=\{D\}\cup\mathcal{D}\setminus\{D_{i}\}$ is again a cut basis. By minimality of $\mathcal{D}$ we must have $|D_{i}|=|D|$ and thus $\mathcal{D}^{\prime}$ is also a minimum cut basis. It follows that $D$ is relevant.
@@
The cuts induced by the edges of a Gomory Hu tree are nested and form a basis of the cut space. The cut-set $\partial A$ of any cut $(A,\bar{A})$ of $G=(V,E,w)$ therefore can be expressed as a linear combination of the cut sets defined by a Gomory-Hu tree $T^{*}$. More precisely,
$\partial A=\bigoplus_{\begin{subarray}{c}uv=e\in F\\ u\in A,v\in\bar{A}\end{subarray}}\partial V_{e}$ (2)
where $\partial V_{e}$ denotes the cutset of the cut $(V_{e},\bar{V}_{e})$ obtained by removing $e$ from $T^{*}$. In other words, $\partial A$ is the $\oplus$-sum of the cuts defined by edges in the Gomory-Hu tree $T^{*}$ with one vertex in $A$ and the other one in $\bar{A}$. Thus $u$ and $v$ are both in $A$ or both in $\bar{A}$ if and only if they are separated by an even number of cut edges along the unique path $P_{uv}$ in $T^{*}$.
@@
A cut $\partial U$ in $G$ is relevant if and only if it is a weight-minimal $u,v$-cut for some pair of distinct vertices $u$ and $v$.
@@
Claim: Every weight-minimal $u,v$-cut is relevant.
@@
Let $D^{*}_{uv}$ be an arbitrary but fixed weight-minimal $u,v$-cut. Sort the list of all cuts in $G$ by non-decreasing weight such that $D^{*}_{uv}$ is the first in the list with its weight. Thus no shorter cut is an $u,v$-cut. Running the canonical greedy algorithm until the list entry before $D^{*}_{uv}$ produces a set $\mathcal{B}^{\prime}$ of linearly independent cuts $\partial U_{i}$ such that $u,v\in U_{i}$ and no $u,v$-cut can be generated. Therefore $\mathcal{B}^{\prime}\cup\{D^{*}_{uv}\}$ is again linearly independent, i.e., the canonical greedy algorithm includes $D^{*}_{uv}$ in the growing basis and hence $D^{*}_{uv}$ is relevant. $\triangleleft$
@@
Claim: If $A$ is not a minimum weight cut for any pair of vertices $u$ and $v$, then it is not relevant.
@@
Let $T^{*}$ be an arbitrary Gomory-Hu tree for $G$. According to equ.(2), $A$ is a linear combination of cut sets $\partial V_{e}$ for an arbitrary Gomory Hu tree $T^{*}$ for the edges that separate vertice $u\in V_{e}$ and $v\in\bar{V}_{e}$. By (G2) and (G3), the weight $\lambda_{e}$ of the edge $e=uv$ in $T^{*}$ satisfies $\lambda_{e}=|\partial V_{e}|$ and, by construction, is the minimum weight of cuts separating $u$ and $v$. Since $A$ is not a minimum weight cut for any pair of vertices, it is in particular not a minimum weight cut for $u$ and $v$. Thus we have $|\partial V_{e}|<|A|$ for every edge $e$ of the $T^{*}$, i.e., $D$ is a linear combination of strictly shorter cuts. By Lemma 3.2, therefore, $A$ it is not relevant. $\triangleleft$
@@
Taken together a cut $D\in\mathfrak{D}$ is relevant if and only if its is a minimum weight cut for at least one pair of vertices $u,v\in V$.
@@
There are graphs with an exponential number of minimal $s,t$-cuts. For instance the $K_{2,n-2}$ with unit edge-weights and $\{s,t\}$ for one part of the bipartition [19]. Since all these cuts are relevant by Thm. 3.6, the number of relevant cuts, like the number of relevant cycles [27], may also grow exponentially with the size of the graph $G$. In the unweighted case, i.e., $w_{pq}=1$ for all edges $pq\in E$, the number of relevant cut is bounded by $n^{\Delta}$, where $\Delta$ is the maximum degree because every minimal $s,t$-cut contains at most $\Delta$ edges and there are no more than $O(n^{\Delta})$ distinct edge sets with at most $\Delta$ edges.
@@
This argument does not generalize to arbitrary weights, however: Consider the graph $G$ obtained from two copies of binary trees with $n$ leaves by identifying corresponding leaves. Thus $G$ has $23n-$ 2 vertices and a maximum vertex degree of $3$. Suppose all edge weights are unity except for edges incident with the leaves, which have weight $1/(2n)$. Denoting the roots of the trees by $s$ and $t$, every cut that contains an edge not incident with a tree-leaf has weight $>1$, while cutting exactly one of the two edges incident with a leaf yields an $s,t$-cut with weight $1/2$. There are $2^{n}$ distinct cuts of this type, all of which a minimum weight $s,t$-cuts.

### v3

@@ Abstract
The set of relevant cuts in a graph is the union of all minimum weight bases of the cut space. We show that a cut is relevant if and only if it is a minimum weight cut between two distinct vertices. Moreover, we give a characterization in terms of Picard-Queyranne Directed Acyclic Graphs that can be used to accelerate the enumeration of the relevant cuts. Finally, we perform an experimental evaluation by comparing with state-of-the-art algorithms.
@@ A Short Note on Relevant Cuts
Nico DomschkeBioinformatics Group, Department of Computer Science, Leipzig University, Härtelstraße 16–18, D-04107 Leipzig, Germanydnico@bioinf.uni-leipzig.de
@@
Richard GolnikBioinformatics Group, Department of Computer Science, Leipzig University, Härtelstraße 16–18, D-04107 Leipzig, Germanyrichard@bioinf.uni-leipzig.de
@@ 1 Introduction
Cuts in graphs are closely related to cycles. The cycle space and the cut space form orthogonal subspaces of the $|E|$-dimensional vector space over GF(2) spanning the edges of a graph $G=(V,E)$. Nevertheless, graphs can contain edge sets called bi-cycles that are both cycles and cuts [27]. The cycle space, and in particular bases of the cycle space comprising short cycles, play an important role in cheminformatics [8, 17, 20] and structural biology [19, 18]. In particular, minimum cycle bases composed of cycles with minimal total length or edge weight, and their union, known as the set of relevant cycles [30], are of use [3].
@@
Minimum cut bases in undirected graphs are studied in detail in [5], their counterparts in directed graphs were analyzed in [15, 16], highlighting the special role of Gomory-Hu trees. These are (not necessarily spanning) trees on $V(G)$ with edge weights $\lambda_{e}$ such that for each pair $s,t\in V(G)$, the minimum weight $\lambda_{e}$ along the path between $s$ and $t$ in $T$ equals the weight of a minimum $s,t$-cut [11]. Since Gomory-Hu trees exist for all graphs, certain sets of n1n-1 distinct minimal cuts already provide a minimal cut for any pair of distinct vertices. As shown in [5, 15], the cuts defined by a Gomory-Hu tree form minimal cut (co-cycle) bases. Moreover, they are weakly fundamental, i.e., their cuts can be ordered in such a way that every cut contains an edge that is not present in any of its predecessors.
@@
Beyond the interest in relevant cuts implied by the analogy to relevant cycles, these are also of practical interest, e.g., for the design of well-behaved crossover operators in graph-based genetic algorithms. For example, recently, cut-and-join crossover operators for chemical graphs have been investigated and benchmarked in detail for search tasks in “chemical space”, demonstrating their efficacy [7].
In the following sections we show that relevant cuts coincide with the cuts that are weight-minimal for some pair of vertices; these results follows from previously known properties of minimal cut bases (see Sections 2 and 3). Then, we use these insights to construct novel algorithms for computing a cut basis and the set of relevant cuts of arbitrary graphs (Section 4). Moreover, we benchmark pre-existing and novel algorithms on chemical and random graphs (Section 5). Finally, we draw our conclusions (Section 6).
@@ 2 Preliminaries
Throughout, $G=(V,E,w)$ is an undirected connected graph with edge weights $w:E\to\mathbb{R}$. In the context of chemical applications, integer weights corresponding to the order of a chemical bond will be most relevant. A cut is a proper bipartition $(V^{\prime},V^{\prime\prime})$ of the vertex set $V(G)$, i.e., $^{\prime}\neq\emptyset, V′′V^{\prime\prime}\neq\emptyset$, $V^{\prime}\cap V^{\prime\prime}=\emptyset$, and $V^{\prime}\cup V^{\prime\prime}=V$. A cutset $D\coloneqq\operatorname{cut}(V^{\prime},V^{\prime\prime})=\{xy\in E|x\in V^{\prime},y\in V^{\prime\prime}\}$. The weight of a cutset is $|D|\coloneqq\sum_{xy\in D}w(xy)$. In the following we will usually write $\partial V^{\prime}=\partial V^{\prime\prime}=\operatorname{cut}(V^{\prime},V^{\prime\prime})=\operatorname{cut}(V^{\prime\prime},V^{\prime})$. In particular, the edges incident to each vertex $x\in V$ form a cutset $\partial\{x\}$. A cutset $D$ that does not include another cutset $D^{\prime}$ of $G$ as a proper subset is called a bond of $G$. Clearly, each $\partial\{x\}$ is a bond. Let $T$ be a spanning tree of $G$, and $e$ an edge of $T$. Removal of $e$ decomposes $T$ into exactly two connected components with vertex sets $V_{e}^{\prime}$ and $V_{e}^{\prime\prime}$. By construction the induced subgraphs $G[V_{e}^{\prime}]$ and $G[V_{e}^{\prime\prime}]$ are connected. The cut $\partial V_{e}^{\prime}$ is also a bond.
@@
for any $W\subseteq V$. For connected graphs, $\mathfrak{D}$ has dimension n1n-1, where $n=|V|$ is the number of vertices in $G$. In particular, each of the sets $\mathcal{D}_{x_{0}}\coloneqq\{\partial\{x\}|x\in V\setminus\{x_{0}\}\}$ for any $x_{0}\in V$, and each of the sets $\mathcal{D}_{T}\coloneqq\{\partial V_{e}^{\prime}|e\in E(T)\}$ for any spanning tree $T$ of $G$ is basis of the cut space, see e.g. [4].
A cut $(V^{\prime},\bar{V}^{\prime})$ is an $s,t$-cut if $s\in V^{\prime}$ and $t\in\bar{V}^{\prime}$. Clearly, the cut space $\mathfrak{D}$ contains an $s,t$-cut for every pair $s\neq t$ of distinct vertices. Less obviously, this is also true for every basis $\mathcal{D}$ of $\mathfrak{D}$, see, e.g., [5]. For completeness, we include a short proof of this statement, which will be used repeatedly throughout this contribution:
@@
Let $\mathcal{D}=\{\partial U_{i}|i=1,\dots,n-1\}$ be a cut-basis comprising cuts $(U_{i},V\setminus U_{i})$ and assume that $\mathcal{D}$ does not contain an $s,t$-cut. Without loosing generality, we may therefore assume that $s,t\in U_{i}$ for all $i$. Consider the set $Z\coloneqq\bigcap_{i}U_{i}$. We have $s,t\in Z$, and moreover, any cut that can be obtained as an $\oplus$-sum of cuts in $\mathcal{D}$ if of the form $(V^{\prime},\bar{V^{\prime}})$ with $Z\subseteq V^{\prime}$ or $Z\subseteq\bar{V^{\prime}}$ since none of the cuts $\partial U_{i}$ contains an edge with both endpoints in $Z$. Hence, note of the cuts $(V^{\prime},\bar{V}^{\prime})$ is an $s,t$-cut, i.e., the $\oplus$-operation on $\mathcal{D}$ cannot generate any $s,t$-cut. The cut space, however, contains in particular the $s,t$-cut $\partial\{s\}$, contradicting our assumption that the $\mathcal{D}$ is a cut basis.
@@
    for every $uv\coloneqq e\in F$, the bipartition of $V$ induced by removing $e$ from $F$ defines a cut of $G$, which we denote by $(V_{e},\bar{V}_{e})$;
@@
    for every $uv\coloneqq e\in F$, the weight $\lambda_{e}$ of $e$ is the weight of a minimum $s,t$-cut;
@@
    for all $s,t\in V$, the weight of a minimum $s,t$-cut in $G$ equals the minimum weight $\lambda_{e}$ along the unique path $P_{st}$ connecting $s$ and $t$ in $T^{*}$.
@@
[11] Every weighted graph $G=(V,E,w)$ has a Gomory Hu tree $T^{*}=(V,F,\lambda)$.
@@
[5] The cutsets $\partial{V_{e}}$ induced by the edges $e\in F$ of a Gomory-Hu tree $T^{*}$ for $G$ form a minimum cut basis of $G$.
While there may be exponentially many minimal $s,t$-cuts for any such fixed node pair, Picard and Queyranne [22] introduced a directed acyclic graph (as a shorthand, we write a PQ-DAG $Q_{s,t}$) that acts as a compact representation and may nevertheless be derived in polynomial time. It is obtained from $G$ by contracting each connected component of the residual flow network to a single node for an arbitrary maximum $s,t$-flow. Accordingly, by construction, vertices in $Q_{s,t}$ represent equivalence classes of vertices w.r.t. the strong connected relationship in the residual flow network, hence sets of nodes from $G$. The “closed” node sets in the DAG $Q_{s,t}$, i.e., those without out-edges, correspond to the minimal $s,t$-cuts.
As shown in [12], given PQ-DAG representations $Q_{u,v}$, corresponding to all minimal $u,v$-cuts, for the edges of the Gomory-Hu tree $uv\in T^{*}$, any PQ-DAG $Q_{s,t}$, where $st$ is not an edge in $T^{*}$, can be computed as a recombination of PQ-DAGs corresponding to edges along the unique path $P_{st}$ connecting $s$ and $t$ in $T^{*}$, improving algorithmic complexity compared to direct construction.
@@ 3 Relevant Cuts
The notion of relevant cycles was studied for cycles bases of undirected graphs in [30]. Interestingly, the notion was proposed already in 1971 by Morris Plotkin in the context of chemical ring perception [23]. More generally, the concept of relevant elements pertains to matroids with a weight function in general and reads as follows for the special case of cut spaces:
@@
The following simple characterization of relevant elements holds for matroids in general [30]:
@@
Since $\mathfrak{D}$ is a vector space it is also matroid. Suppose $D$ is contained in the minimum cut basis $\mathcal{D}$. If $D$ can be written as the $\oplus$-sum of shorter cuts $D_{1}$ through $D_{k}$, it could be exchanged with one of these cuts, contradicting minimality of $\mathcal{D}$. This is justified as follows: We know that at least one cut $D_{1}$ through $D_{k}$ is not element of $\mathfrak{D}$, otherwise $D\notin\mathfrak{D}$ by linear independence. If exactly one cut $D_{i}$ is not an element of $\mathfrak{D}$, $D_{i}$ and $D$ can be exchanged. If multiple cuts are not in $\mathfrak{D}$, we know they can be constructed as $\oplus$-sums of yet shorter cuts, and we apply this argument recursively.
Now suppose that $D$ cannot be written as a sum of shorter cuts. Then $D=D_{1}\oplus\dots\oplus D_{k}$ and there is at least one cut $D_{i}$ with $|D_{i}|\geq|D|$. Now $D_{i}=D_{1}\oplus\dots\oplus D_{i-1}\oplus D\oplus D_{i+1}\oplus\dots\oplus D_{k}$ and thus $\mathcal{D}^{\prime}=\{D\}\cup\mathcal{D}\setminus\{D_{i}\}$ is again a cut basis. By minimality of $\mathcal{D}$ we must have $|D_{i}|=|D|$ and thus $\mathcal{D}^{\prime}$ is also a minimum cut basis. It follows that $D$ is relevant.
@@
The cuts induced by the edges of a Gomory Hu tree are nested and form a basis of the cut space. The cutset $\partial V^{\prime}$ of any cut $(V^{\prime},\bar{V^{\prime}})$ of $G=(V,E,w)$ therefore can be expressed as a linear combination of the cut sets defined by a Gomory-Hu tree $T^{*}$. More precisely,
$\partial V^{\prime}=\bigoplus_{\begin{subarray}{c}uv=e\in F\\ u\in V^{\prime},v\in\bar{V^{\prime}}\end{subarray}}\partial V_{e}$ (2)
where $\partial V_{e}$ denotes the cutset of the cut $(V_{e},\bar{V}_{e})$ obtained by removing $e$ from $T^{*}$. In other words, $\partial V^{\prime}$ is the $\oplus$-sum of the cuts defined by edges in the Gomory-Hu tree $T^{*}$ with one vertex in $V^{\prime}$ and the other one in $\bar{V^{\prime}}$. Thus $s$ and $t$ are both in $V^{\prime}$ or both in $\bar{V^{\prime}}$ if and only if they are separated by an even number of cut edges along the unique path $P_{st}$ in $T^{*}$.
@@
A cut $\partial V^{\prime}$ in $G$ is relevant if and only if it is a weight-minimal $s,t$-cut for some pair of distinct vertices $s$ and $t$.
@@
Claim: Every weight-minimal $s,t$-cut is relevant.
@@
Let $D^{*}_{st}$ be an arbitrary but fixed weight-minimal $s,t$-cut. Sort the list of all cuts in $G$ by non-decreasing weight such that $D^{*}_{st}$ is the first in the list with its weight. Thus, no shorter cut is an $s,t$-cut. Using the canonical greedy algorithm [9], a minimal cut basis $\mathcal{B}$ is obtained by iterating over the sorted lists of cuts, adding a cut $D$ to the initial empty set $\mathcal{B}$ if and only if $\mathcal{B}\cup\{D\}$ is linearly independent. Now consider the set $\mathcal{B}^{\prime}$, created by the above procedure up to the point where all cuts have been processed that appear in the list before the first $s,t$-cut $D^{*}_{st}$. By construction, $\mathcal{B}^{\prime}$ is linearly independent and each cutset $\partial U_{i}\in\mathcal{B}^{\prime}$ corresponds to a cut $(U_{i},\bar{U}_{i})$, where we may assume without loosing generality that $s,t\in U_{i}$. By the same argument as in the proof of Lemma 2.1, $\mathcal{B}^{\prime}$ cannot generate an $s,t$-cut as an $\oplus$-sum. Thus $\mathcal{B}^{\prime}\cup\{D^{*}_{st}\}$ is again linearly independent. In the next step, the canonical greedy algorithm therefore includes $D^{*}_{st}$ in the growing minimal weight basis $\mathcal{B}^{\prime}$. It follows that $D^{*}_{st}$ is relevant. $\triangleleft$
@@
Claim: If $\partial V^{\prime}$ is not a minimum weight cut for any pair of vertices $s$ and $t$, then it is not relevant.
@@
Let $T^{*}$ be an arbitrary Gomory-Hu tree for $G$. According to Eq. (2), $\partial V^{\prime}$ is a linear combination of cutsets $\partial V_{e}$ defined by $T^{*}$, where the edges $e$ separate vertices $s\in V_{e}$ and $t\in\bar{V}_{e}$. By (G2) and (G3), the weight $\lambda_{e}$ of the edge $e$ in $T^{*}$ satisfies $\lambda_{e}=|\partial V_{e}|$ and, by construction, is the minimum weight of cuts separating $s$ and $t$ if $\lambda_{e}$ is the minimal weight encountered along the unique path $P_{st}$ in $T^{*}$. Since $\partial V^{\prime}$ is not a minimum weight cut for any pair of vertices, it is in particular not a minimum weight cut for $s$ and $t$. Thus we have $|\partial V_{e}|<|\partial V^{\prime}|$ for every edge $e$ of the $T^{*}$ appearing in Eq. (2). Thus $D$ is a linear combination of strictly shorter cuts. By Lemma 3.2, therefore, $\partial V^{\prime}$ it is not relevant. $\triangleleft$
@@
Taken together, a cut $D\in\mathfrak{D}$ is relevant if and only if it is a minimum weight cut for at least one pair of vertices $s,t\in V$.
@@
There are graphs with an exponential number of minimal $s,t$-cuts. For instance the $K_{2,n-2}$ with unit edge-weights and $\{s,t\}$ for one part of the bipartition [22]. Since all these cuts are relevant by Thm. 3.6, the number of relevant cuts, like the number of relevant cycles [30], may also grow exponentially with the size of the graph $G$. In the unweighted case, i.e., $w_{pq}=1$ for all edges $pq\in E$, the number of relevant cuts is bounded by $|E|^{\Delta}$, where $\Delta$ is the maximum degree because every minimal $s,t$-cut contains at most $\Delta$ edges and there are no more than $O(|E|^{\Delta})$ distinct edge sets with at most $\Delta$ edges.
@@
This argument does not generalize to arbitrary weights, however: Consider the graph $G$ obtained from two (potentially identical) instances of binary trees with $n$ leaves by identifying corresponding leaves. Thus $G$ has 3n23n-2 vertices and a maximum vertex degree of 33. Suppose all edge weights are unity except for edges incident with the leaves, which have weight $1/(2n)$. Denoting the roots of the trees by $s$ and $t$, every cut that contains an edge not incident with a tree-leaf has weight >1>1, while cutting exactly one of the two edges incident with a leaf yields an $s,t$-cut with weight $1/2$. There are $2^{n}$ distinct cuts of this type, all of which a minimum weight $s,t$-cut (see Fig. 1).
st1110.10.10.10.10.10.10.10.10.1110.11
Figure 1: Example of a graph with arbitrary weights and an exponential number of $s,t$-cuts. Following the construction in the main text joining two binary trees (solid vs dashed lines) with $n=5$ leaves, one edge adjacent per leave node has to be included in each minimal cut, yielding $2^{5}$ possibilities. Two such minimal cuts are marked in blue and red.